### v1

@@ 4.2 Opportunities for future work
This study presents several avenues for future research, each addressing current limitations and opportunities for expansion. One primary constraint lies in the data quality of OpenAlex, which has been shown to have missing information issues, particularly in author affiliation and abstract details [32]. These gaps could potentially lead to a loss of valuable insights from papers or preprints not fully captured by OpenAlex. Future research could benefit from conducting sensitivity analyses to assess the impact of these missing values on the patterns observed in this study, thereby providing a more robust understanding of the data’s reliability and comprehensiveness.
@@
Last, while this study conducts a before-and-after comparison, the observed differences should not be interpreted as causal, given the inherent time-varying trends in collaboration along with other confounding effects are not well controlled. Preliminary analyses using the Bayesian Structural Time Series (BSTS) model are conducted in this study (see Figure 7); however, a more robust quasi-experimental design incorporating control and treatment groups could merit further exploration to infer the causal effects of LLMs on collaboration patterns.
@@ A.1 Data preparation workflow
Figure 4 outlines the process for collecting and cleaning the dataset of LLM-relevant papers for our analysis. The process begins with a broad search using general terms related to LLMs and popular models based on the MMLU benchmark (see subsection A.2) [24], spanning from October 2018 to June 2024. We apply this search to the title and abstract to avoid excessive noise in the dataset. This initial search yields a collection of 97,242 papers, which then undergo a series of filtering steps to enhance relevance and remove duplicates.
The cleaning process involves several steps, each progressively narrowing down the dataset. First, the collection is limited to preprints and articles, reducing it to 87,489 papers. The next step involves removing duplicates based on titles. Additionally, since a preprint might change its title upon official publication, we examine papers with slight variations between preprints and formal publications using Jaccard similarity (see subsection A.3). After reviewing the papers with Jaccard similarity, we find that a paper with a similarity score greater than 0.6 is very likely to be duplicated, and we remove those potential duplicates. However, it is still possible for a paper to contain LLM-related keywords but not relevant to LLM studies, such as the keyword “PaLM” which often appears in papers discussing “palm trees” or “palm oil.” To address this issue, we employ the GPT-4o model to support the evaluation of the relevance of a paper (see subsection A.4). This filtering process results in a final set of 50,391 papers, representing a focused and highly relevant collection for our analysis. The distribution of these 50,391 papers by field is presented in Figure 1.
@@ A.2 Paper collection
Table 2 below shows the specific search terms used to collect papers from OpenAlex. The search terms include two general terms, “large language model” and “LLM,” along with popular open-source models (e.g., BERT, Flan-T5, LLaMA) and closed-source models (e.g., ChatGPT, Claude) based on the MMLU benchmark [24].
@@
Close-sourced (GPT 2) OR (GPT 3) OR (GPT 4) OR ChatGPT OR (Claude Instant) OR (Claude 1.3) OR (Claude 2) OR (Claude 3) OR (Google PaLM) OR (Google Gemini) OR (PaLM 2) OR (Gemini Pro) OR (Gemini Ultra) OR (davinci-002) OR (davinci-003) OR (Chinchilla 70B)
Open-sourced BERT OR RoBERTa OR (Meta LLaMA) OR (LLaMA 2) OR (LLaMA 3) OR Mistral OR Mixtral OR Qwen OR DBRX OR (Falcon 40B) OR (Falcon 7B) OR (Falcon 180B) OR (OPT 66B) OR (BLOOM 176B) OR (GLM 130B) OR (Flan T5) OR (Flan PaLM)
@@ A.3 Duplicates removal
We first remove the duplicated papers based on the title information. This handling reduces the collection of papers to 70,644. Given that a preprint might change its title when officially published, we then use the Jaccard similarity coefficient to identify works with variations in titles but are actually the same. The Jaccard similarity coefficient is a measure of similarity between two sets. It is defined as the size of the intersection divided by the size of the union of the sets, as shown below:
@@ A.4 Relevance evaluation
This step involves filtering out papers that are not actually discussing LLM-related topics, even if their abstracts contain key terms such as “Palm.” We take two steps to do the filtering. First, we consider papers with identified topics (provided by OpenAlex) including “natural language processing,” “artificial intelligence,” “machine learning,” “text mining,” “deep learning,” “transfer learning,” “question answering,” and “speech recognition” as LLM-relevant. Then, we use GPT-4 to determine their relevance based on the title, abstract, and topics. The prompt is designed as follows:
@@
To validate the classification, we manually reviewed 190 randomly selected papers, of which 86 are classified as irrelevant and 104 as relevant based on our manual annotation. We then compare these results with those returned by GPT-4o. We use F1-score and accuracy to evaluate performance. Overall, GPT-4’s identification achieves an agreement of 0.96 with manual annotation, with F1-scores of 0.96 for both irrelevant and relevant classes.
@@ B.1 Wilcoxon rank-sum test
B.2 Bayesian structural time series analysis
The entropy changes after the launch of ChatGPT cannot be solely attributed to causal effects due to inherent trends in entropy over time. To address this, a Bayesian structural time series (BSTS) model is fitted for each field for time series forecasting and causal inference [33, 34]. Figure 7 displays results where the impact of ChatGPT is statistically significant. Trained by pre-intervention data, the BSTS model predicts post-intervention entropy (the dashed line in each panel’s top subfigure). The difference between these predictions and observations provides an estimate of the intervention’s pointwise effect (second subfigure in each panel), and their cumulative sum indicates the total effect over time (third subfigure in each panel). As shown, the release of ChatGPT consistently leads to a statistically significant increase in the entropy in Computer Science. It also leads to a statistically significant decrease in the entropy in Medicine based on institutions, and a statistically significant decrease in the entropy in Business, Management & Accounting based on departments. Other fields, however, do not show significant impacts by ChatGPT based on the BSTS causal inference.
Figure 7: BSTS outcomes for entropy based on authors’ affiliated (a) institutions and (b) departments. The results plotted only include fields where the release of ChatGPT in November 2022 has shown a statistically significant impact.
@@ Appendix C OpenAlex topic and field classification
Appendix D Network analysis for countries
The network metrics based on authors’ affiliated country information are presented in Table 3, and the network visuals are presented in Figure 8.
Figure 8: Co-authorship networks based on countries of researchers. Each node represents a researcher’s country of affiliation, and each edge represents a co-authorship between pairs of researchers from the respective country. Node colors represent the clusters to which the nodes belong, determined based on Louvain modularity. Node labels represent the top 20% of nodes ranked by degree centrality. Edge thickness represents the frequency of co-authorship between the connected researchers.
Table 3: Network metrics based on authors’ affiliated countries
Metric Countries
No. Nodes 155
No. Edges 1745
Density 0.15
Avg. Shortest Path 9.36
Clustering Coeff. 0.77
No. Components 16
Largest Component 140
Power-law Exponent 5.22
No. Comm.(Louvain) 20
No. Comm. (CNM) 19
Degree Centrality (top 5) US (109), GB (91), CN (81), DE (79), CA (75)
Betweenness Centrality (top 5) PL (0.056), GB (0.053), CA (0.051), FR (0.050), TR (0.050)
Closeness Centrality (top 5) PT (0.157), MX (0.157), QA (0.155), TR (0.154), SY (0.154)
Eigenvector Centrality (top 5) US (0.630), CN (0.498), GB (0.395), DE (0.222), CA (0.194)

### v2

@@ 4.2 Opportunities for future work
This study presents several avenues for future research, each addressing current limitations and opportunities for expansion. One primary constraint lies in the data quality of OpenAlex, which has been shown to have missing information issues, particularly in author affiliation and abstract details [49]. To mitigate this, our analysis excludes records with incomplete data. Consequently, approximately 50% of the papers retain complete information on authors’ institutions, and about 40% retain detailed department information. While this approach helps address the data quality issue, it also leads to the exclusion of valuable insights from papers or preprints that are not fully captured by OpenAlex, especially in fields with fewer published papers. Nevertheless, as highlighted in prior research [50], OpenAlex continues to demonstrate better inclusiveness and coverage of academic papers compared to other databases. Future research could benefit from conducting sensitivity analyses to assess the impact of these missing values on the patterns observed in this study, thereby providing a more robust understanding of the data’s reliability and comprehensiveness.
The publication timing of a paper could potentially influence our results. In this study, we use the publication time of a paper as the reference for time. However, for papers without a specific publication date—particularly those from Computer Science conferences, which often only provide the publication year by OpenAlex—we utilize the created_date field from the OpenAlex dataset as a proxy for the publication time. This approach ensures the inclusion of papers from conference tracks while providing a reasonable approximation for their publication dates. Future research could incorporate the information of papers from Computer Science conference tracks available in their respective databases. However, ensuring consistency during the data fusion process presents an additional challenge that would need to be addressed.
@@ A.1 Data preparation workflow
Figure 6 outlines the process for collecting and cleaning the dataset of LLM-relevant papers for our analysis. The process begins with a broad search using general terms related to LLMs and popular models based on the MMLU benchmark (see subsection A.2) [35], spanning from October 2018 to September 2024. We apply this search to the title and abstract to avoid excessive noise in the dataset. This initial search yields a collection of 177,462 papers, which then undergo a series of filtering steps to enhance relevance and remove duplicates.
The cleaning process involves several steps, each progressively narrowing down the dataset. First, the collection is limited to preprints and articles, reducing it to 128,736 papers. The next step involves removing duplicates based on titles. Additionally, since a preprint might change its title upon official publication, we examine papers with slight variations between preprints and formal publications using Jaccard similarity (see subsection A.3). After reviewing the papers with Jaccard similarity, we find that a paper with a similarity score greater than 0.6 is very likely to be duplicated, and we remove those potential duplicates. However, it is still possible for a paper to contain LLM-related keywords but not relevant to LLM studies, such as the keyword “PaLM” which often appears in papers discussing “palm trees” or “palm oil.” To address this issue, we employ the GPT-4o model to support the evaluation of the relevance of a paper (see subsection A.4). This filtering process results in a final set of 59,293 papers, representing a focused and highly relevant collection for our analysis. The distribution of these papers by field is presented in Figure 1.
@@ A.2 Paper collection
Table 3 below shows the specific search terms used to collect papers from OpenAlex. The search terms include two general terms, “large language model” and “LLM,” along with popular open-source models (e.g., BERT, Flan-T5, LLaMA) and closed-source models (e.g., ChatGPT, Claude) based on the MMLU benchmark [35]. It should be noted that our search term like “GPT” encompasses all relevant models, including GPT-3, GPT-4, and GPT-4o.
@@
Close-sourced GPT OR ChatGPT OR Claude OR Gemini OR (PaLM 2) OR (davinci-002) OR (davinci-003) OR Chinchilla
Open-sourced BERT OR RoBERTa OR LLaMA OR Mistral OR Mixtral OR Qwen OR DBRX OR Falcon OR BLOOM OR (Flan T5) OR (Flan PaLM) OR LaMDA
The control groups follow the same process. For the “ML” control group, we randomly select 100,000 papers that containing the keyword “machine learning” from the same time periods and apply the processing pipeline. This process reduces the dataset to 71,482 papers. To ensure that the collected ML papers are distinct from those in the LLM group, we remove any overlapping papers that appear in the LLM group. Similarly, for the “Non-LLM/ML” papers, we initially collect a random sampling of 150,000 papers. Then, we follow the same cleaning procedure and eliminate overlaps. In other words, each group is mutually exclusive, with no papers belonging to more than one group. As a result, we have 59,293, 70,945, and 73,110 papers for “LLM,” “ML,”, and “Non-LLM/ML” groups, respectively.
@@ A.3 Duplicates removal
We first remove the duplicated papers based on the title information. This handling reduces the collection of papers to 128,736. Given that a preprint might change its title when officially published, we then use the Jaccard similarity coefficient to identify works with variations in titles but are actually the same. The Jaccard similarity coefficient is a measure of similarity between two sets. It is defined as the size of the intersection divided by the size of the union of the sets, as shown below:
@@ A.4 Relevance evaluation
This step involves filtering out papers that are not actually discussing LLM-related topics, even if their abstracts contain key terms such as “Palm.” We take several factors into consideration for this filtering process. First, we consider papers with identified topics (provided by OpenAlex) including “natural language processing,” “artificial intelligence,” “machine learning,” “text mining,” “deep learning,” “transfer learning,” “question answering,” and “speech recognition” as LLM-relevant. Then, we use GPT-4o mini to determine their relevance based on the title, abstract, and topics. The prompt is designed as follows:
@@
To validate the classification, we manually review 190 randomly selected papers, of which 86 are classified as irrelevant and 104 as relevant based on our manual annotation. We then compare these results with those returned by GPT-4o mini. We use F1-score and accuracy to evaluate performance. Overall, GPT-4o mini’s identification achieves an agreement of 0.96 with manual annotation, with F1-scores of 0.96 for both irrelevant and relevant classes.
A.5 Discipline-related information extraction
As the department information is not directly provided by OpenAlex, this step involves extracting discipline-related information from the authors’ affiliation information, which allows us to conduct department analysis in the result section. Again, we use GPT-4o mini and design the following prompt to facilitate the information retrieval. For example, if an author’s affiliation information is “School of Computer Science, Peking University, Beijing, China,” then this prompt allows us to extract the information “Computer Science” as the relevant discipline. It is important to note that we do not merge closely related disciplines, such as “Computer Science” and “Computer Engineering,” to ensure precise and unbiased extraction.
Please extract the discipline-related information from the following
authors affiliation:
Authors affiliation: {affiliation}
For example, for Department of Biological Science, Joseph Ayo Babalola
University, Nigeria’, return Biological Science’. If it is not written
in English, translate it to English and return. If you cannot identify
any discipline-related information, return None’.
Please do not return other output or explanation.
@@ Appendix C OpenAlex topic and field classification
Appendix D Top centrality entities
Table 4 and Table 5 present the top five entities based on degree centrality, betweenness centrality, closeness centrality, and eigenvector centrality, calculated from the authors’ institutional and departmental affiliations, respectively. Each table lists the top five entities across three paper groups.
Table 4: Top 5 entities based on authors’ institution affiliations
Metric LLM ML Non-LLM/ML
Degree Centrality Stanford University (342) Stanford University (667) Centre National de la Recherche Scientifique (532)
Tsinghua University (274) University College London (612) University College London (437)
Peking University (263) University of Oxford (561) University of Toronto (393)
University of Washington (261) Harvard University (529) University of California, Los Angeles (393)
Harvard University (245) University of Cambridge (523) University of Oxford (376)
Betweenness Centrality Instituto de Ciencias de La Construcción Eduardo Torroja (0.500) Imperial College London (0.019) King Saud University (0.017)
Escuela Nacional de Sanidad (0.500) University College London (0.017) University of Toronto (0.016)
Heilongjiang Academy of Sciences (0.500) University of Oxford (0.016) University of Oxford (0.014)
China Academy of Space Technology (0.500) Stanford University (0.015) Chinese Academy of Sciences (0.013)
Nahrain University (0.500) New York University (0.014) Centre National de la Recherche Scientifique (0.012)
Closeness Centrality University College London (0.022) Trilogi University (0.500) African Studies Centre (0.500)
University of Waterloo (0.021) Universitas Bina Darma (0.500) Institute for Development & Economic Analysis (0.500)
Yale University (0.017) Nahrain University (0.500) Semmes Murphey Foundation (0.500)
Harvard University (0.017) Heilongjiang Academy (0.500) Kaleida Health (0.500)
Peking University (0.016) China Academy Space Tech. (0.500) Buffalo General Medical Center (0.500)
Eigenvector Centrality Shanghai Jiao Tong University (0.339) Institute of High Energy Physics (0.503) Institute of High Energy Physics (0.521)
Shanghai Sixth People’s Hospital (0.316) Université Libre de Bruxelles (0.470) Université Libre de Bruxelles (0.453)
Stanford University (0.276) UCLouvain (0.415) Vrije Universiteit Brussel (0.402)
Tsinghua University (0.254) Vrije Universiteit Brussel (0.370) UCLouvain (0.372)
Chinese University of Hong Kong (0.252) Ghent University Hospital (0.319) University of Antwerp (0.329)
Table 5: Top 5 entities based on authors’ department affiliations
Metric LLM ML Non-LLM
Degree Centrality Computer Science (1392) Computer Science (3001) Medicine (1948)
Medicine (1019) Medicine (2184) Chemistry (1278)
Computer Science and Engineering (611) Engineering (1501) Physics (1007)
Artificial Intelligence (586) Physics (1240) Public Health (996)
Engineering (385) Computer Science and Engineering (1201) Biology (879)
Betweenness Centrality Dynamics, Logics and Inference for Biological Systems (0.500) Computer Science (0.0761) Medicine (0.041)
Data and Knowledge Management (0.500) Medicine (0.0372) Chemistry (0.037)
Computer Engineering and Interdisciplinary Major (0.500) Engineering (0.0429) Mechanical Engineering (0.030)
Control & Automation Engineering (0.500) Mechanical Eng. (0.0298) Engineering (0.029)
Reasoning (0.500) Physics (0.0273) Physics (0.029)
Closeness Centrality Computer Science (0.091) Innovation & Society (0.500) Rail Systems Management (0.500)
Medicine (0.043) Responsible Innovation (0.500) Aviation Management (0.500)
Artificial Intelligence (0.036) Data & Knowledge Mgmt. (0.500) Computing Science (0.500)
Computer Science and Engineering (0.033) Maritime Transportation Safety (0.500) Control, Robotics, and Electrical Eng. (0.500)
Education (0.031) Political Science and International Relations (0.500) Applied Analysis and Stochastics (0.500)
Eigenvector Centrality Business (0.561) Physics (0.7022) Physics (0.698)
Information Systems (0.311) Astronomy (0.102) Astronomy (0.102)
Technology, Policy and Management (0.070) Astrophysics (0.056) Nuclear Physics (0.090)
Management Studies (0.053) Nuclear Physics (0.053) Astrophysics (0.056)
Marketing and Marketing Management (0.053) Earth Sciences (0.022) Astroparticle and Cosmology (0.050)